OpenStudy (anonymous):

Laura owes $4759 on a credit card with a 20.9% interest rate compounded monthly. What is the monthly payment she should make in order to pay off this debt in 3 years, assuming she does not charge any more purchases with the card? (I just need the formula I need to use for this)

OpenStudy (wolf1728):

If the formula you need is the loan payment formula, then try this link: http://www.1728.org/loanform.htm I used a calculator located here: http://www.1728.org/calcloan.htm to compute the monthly payment as $179.05 but we can work this out with the formula.

OpenStudy (wolf1728):

To use the formula we first must divide the rate 20.9 by 1200 which equals 0.01741667 Monthly Payment = rate + (rate/[(1+rate)^months)-1] * principal Monthly Payment = 0.01741667 + (0.01741667/([(1.01741667)^36]-1) * principal Monthly Payment = 0.01741667 + 0.01741667 / [(1.8619095005)-1] * principal Monthly Payment = 0.01741667 + (0.01741667 / 0.8619095005) * principal Monthly Payment = 0.01741667 + 0.0202070751 * principal Monthly Payment = 0.0376237451 * principal Monthly Payment = 0.0376237451 * 4,759 Monthly Payment = 179.0514028977 = $179.05 See ? It's just that simple!! LOL